### v2

# Introducing AI to an Online Petition Platform Changed Outputs but not Outcomes

Abstract
The rapid integration of AI writing tools into online platforms raises critical questions about their impact on content production and outcomes. We leverage a unique natural experiment on Change.org, a leading social advocacy platform, to causally investigate the effects of an in-platform “write with AI” tool. To understand the impact of the AI integration, we collected 1.5 million petitions and employed a difference-in-differences analysis. Our findings reveal that in-platform AI access significantly altered the lexical features of petitions and increased petition homogeneity, but did not improve petition outcomes. We confirmed the results in a separate analysis of repeat petition writers who wrote petitions before and after introduction of the AI tool. The results suggest that while AI writing tools can profoundly reshape online content, their practical utility for improving desired outcomes may be less beneficial than anticipated, and introduce unintended consequences like content homogenization.
01_introduction
Results
To estimate the causal impact of access to the Change.org in-platform AI tool, we conducted a series of difference-in-difference analyses [callaway2021difference] that leveraged a natural experiment in AI tool access on Change.org driven by a staggered roll-out by geographic region: the United States, Great Britain, and Canada received full access ten weeks prior to Australia. [^1] We confirmed country-specific launch dates for the AI-generated draft tool with the Change.org team. We could thus compare country-level trends in petition lexical style and outcomes during the ten-week period where access differed between geographic regions.
The hypothesis underlying the analyses is that since all countries followed similar trends in petition style and outcomes prior to AI tool release, we can attribute differences in trends during the ten-week period to the introduction of the AI tool. We verify the existence of similar trends through a parallel trends analysis and time series figures (included in the SI Appendix) in which we show that before the release of AI, and after all countries receive AI access, trends in style and outcomes of petitions produced in the control country (Australia) were parallel to those in treated countries (the United States, Great Britain, Canada). Our comparison of trends followed a difference-in-differences quasi-experimental approach, in which we measured the change in trends between similar groups after some groups received an intervention. We focus on the ten weeks of differential access between the launch of the AI tool in treated countries and the launch of the AI tool in Australia. We leverage two types of specifications for the difference-in-differences approach: static and dynamic. In the static difference-in-differences specification, we measure the overall effect across countries and weeks of AI access for treated countries. In the dynamic difference-in-differences specification, we measure the effect of AI access for each week in which treated countries have AI access [callaway2021difference]. We also confirm our results through a series of robustness tests and a synthetic controls analysis, which reaffirm our difference-in-differences results (see SI Appendix).
We defined the intervention in our causal analysis as AI access, rather than the actual use of AI in the petition. While we know when users had access to the in-platform AI tool, we cannot robustly tell when and how they used it, as AI detection models may be biased in their ability to identify AI-generated content. Users may also use off-platform generic AI tools; however, we expect off-platform usage to be similar in both treated and control countries and thereby to not influence our estimates. We did develop an AI classifier specific to the Change.org AI tool to understand if the tool was utilized when users had access. With this method, we estimate that during the ten-week period of differential access, over half of petitions that were written with access to the in-platform AI tool had used the tool (see SI Appendix). We also find further support for parallel trends: Australia and treated countries had similar detected levels of (general, off-platform) AI use prior to the release of the in-platform AI tool.
In our main analysis we use Australia as the control country to estimate causal effects. We confirm that our estimates are robust to multiple geographic regions with a synthetic control analysis that leverages additional control countries (New Zealand, India, and China). The results of the synthetic control analysis are consistent with the difference-in-differences analysis, demonstrating that our causal estimates are consistent across regions and outside of Australia.
We preregistered our difference-in-differences analysis with AsPredicted, and note below the few instances where we expanded from the preregistration.
Figure 1: Platform trends in lexical features (a, b) changed significantly with AI access while trends in outcomes (c, d) did not improve, or worsened. Difference-in-differences (DiD) analysis of platform trends in petitions’ lexical features (a, b) and outcomes (c,d) before and after Change.org AI tool release. Change in outcomes reflect change in percentage points. (a) Dynamic difference-in-differences with estimated effect of AI access on lexical features by week, relative to the baseline week (Week -1). (b) Aggregated estimated effect of AI access on lexical features, relative to pre-AI period. (c) Dynamic difference-in-differences with estimated effect of AI access on outcomes by week. (d) Aggregated estimated effect of AI access on outcomes, relative to pre-AI period. The A/B testing period, shown in gray between the dashed lines, is excluded from analysis. 95% confidence intervals are shown. Pre-AI period was 64 weeks (N = 235,947), A/B test period was 27 weeks (N = 82,781), and the Post-AI period was 10 weeks (N = 42,100).
Lexical Features
We started by estimating the effect of in-platform AI tool access on petition lexical style through three features: lexical diversity, readability, and text length. We selected these lexical features because of their strong predictive relationship to petition outcomes prior to the release of the AI tool (see SI Appendix). Lexical diversity and readability have been used to assess writing competency in a variety of contexts such as language proficiency and clarity [mesgar2018neural, bizzoni2023good], and sufficient length is necessary for a petition to convey a compelling message. We operationalized lexical diversity through Moving Average Type-Token Ratio (MATTR), a measure of lexical diversity calculated as a moving average of the ratio of unique words to total words across windows of the text. Higher MATTR scores denote a more varied vocabulary. We operationalized readability using the Flesch-Kincaid Grade Level (FKGL), a common measure of readability calculated with the number of words per sentence and syllables per word. Higher FKGLs denote more complex text that requires a higher grade level to understand. For text length we used a simple word count of the number of words in the petition text and title combined.
Figure 1(a) demonstrates that access to the AI tool had a significant and consistent effect on lexical features: petitions were longer, with more complicated and varied vocabularies. Figure 1(a) shows the results of a dynamic difference-in-differences estimation, in which we estimate the average treatment effect in treated countries by week (x-axis), compared to the baseline (week -1). The figure shows average treatment effects for treated countries in weeks before any AI tool access (orange, left of week 0), during AI tool A/B testing (gray, weeks 0-25), and after the full launch of the tool (blue, right of week 25). The shift (y-axis) is captured via average group-time treatment effect estimates on the treated users, relative to the baseline week (week -1). [^2] We exclude the A/B period from our analysis (but show it in the figure) as we do not know which users or what share of users accessed the tool during this period. The spikes in estimated effect on petition style during the A/B test period reflect experimentation with prompt engineering that informed the in-platform AI tool. After full launch, the model and prompt behind the AI tool remain consistent during the ten-week period for which we estimate the effect of AI tool access. For example, the Word Count graph in Figure 1(a) shows that in weeks after full launch (blue, right), AI access resulted in an estimated addition of 50-70 words to the median petition word count, relative to the baseline week. Similarly, access to the AI tool increased median MATTR values and FKGLs in weeks with AI access.
Figure 1(b) shows the result of a static difference-in-differences estimation, by which the average treatment effect is estimated across the post-AI (after week 25) period, relative to the pre-AI period (before week 0). We estimate that over all petitions in the post-AI period, the MATTR (top row of Figure 1(b)) increased by 0.049 units (95% CI = [0.048, 0.050]), median petition grade level (second row of Figure 1(b)) increased by 1.51 grades (95% CI = [1.22, 1.79]), and median word count (bottom row of Figure 1(b)) increased by 55.15 words (95% CI = [48.63, 61.68]). An estimated increase of 50 words corresponds to a 45% increase in median petition length, from 110 words pre-AI, during this ten-week period. Similarly, the median FKGL increases by 19%, from grade 8 pre-AI, and the MATTR increases by 6%, from 0.8 units pre-AI. A descriptive analysis shows that petition vocabularies qualitatively changed as well: prior to AI access the most frequently used verbs in petition titles were mono-syllabic Old English verbs such as “stop” and “let”. Following the release of the AI tool, multi-syllabic verbs like “implement” and “establish” jumped in frequency (see time series of verb frequency in the SI Appendix).
Figure 2: Dynamic difference-in-difference estimates for homogeneity in countries with access to AI shows that with access to the AI tool, platform homogeneity increases. Estimated average treatment effect (ATT) of mean pairwise similarity is shown with 95% CIs. Estimates are compared to the baseline the week prior to A/B testing in treated countries (Week -1). Pre-AI period was 64 weeks (N = 235,947), A/B test period was 27 weeks (N = 82,781), and the Post-AI period was 10 weeks (N = 42,100).
Platform Homogeneity
We augmented our pre-registered difference-in-differences analysis by studying the impact of AI access on platform homogeneity. The results show an increase in inter-petition homogeneity, as measured through the mean pairwise similarity between petitions [wu2024generativemonoculturelargelanguage, padmakumar2023does]. To measure homogeneity, we first created 3,072-dimension embeddings with the OpenAI text-embedding-3-large model [^3] https://platform.openai.com/docs/models/text-embedding-3-large for all petitions. We then calculated the cosine similarity between all pairs of petitions posted in a given week and country, and took the mean of these pairwise scores to obtain the mean pairwise similarity for that week and country. We used these weekly country-level time series of homogeneity to compare trends in treated and untreated countries through a dynamic difference-in-differences estimation.
The dynamic difference-in-differences results shown in Figure 2 demonstrate that access to the AI tool caused petitions to be more similar to one another than if they had been written without access to the AI tool. In weeks after week 25 (shown in blue), following the introduction of AI, AI access had a consistent effect on petition homogeneity, increasing mean pairwise similarity by about 0.05 units (on a scale of -1 to 1) for each period. We also conducted a static difference-in-difference estimation (shown in the SI Appendix), finding that overall, the post-AI period had a 0.055 (95% CI = [0.051, 0.059]) increase in pairwise similarity, which is a 9% increase from the average pre-AI weekly mean similarity value of 0.583.
Outcomes
We find that petitions’ outcomes did not improve—and in some cases even worsened—with access to the AI tool. We examined petition outcomes as the share of petitions that reach 1 comment in 30 days and the share that reach a total of 10 signatures at the time of data collection. We preregistered our primary outcome as the presence of a comment posted within 30 days after the petition was posted, since the time of comments (but not signatures) is available in our data. As each petition was posted at least one year prior to data collection, we expect recency was not a factor in whether petitions reached these minimal thresholds. For example, 99% of petitions receive their first comment in the first month after posting, if they receive a comment at all. We include additional analysis with different thresholds for comments and signatures in the SI appendix, showing that results remain consistent. We use these minimal thresholds instead of raw engagement counts because engagement is sparse (only 28.17% of petitions reach 10 signatures and 26.65% ever gain a single comment). Furthermore, petitions on Change.org concern a wide variety of communities, from hyperlocal neighborhood issues that persuade decision-makers with tens of supporters, to global issues that gain hundreds of thousands of signatures. On the other hand, the minimum thresholds for petitions to “get off the ground” remain the same across time and context. By using minimum thresholds we ensure our analysis is not biased towards issues that affect more populous communities while still reflecting a meaningful indicator of a petition’s outcomes.
We do not find evidence that AI tool access positively impacted petition outcomes. Instead, we observe that with AI access, outcomes are either unchanged, with no significant differences observed post-treatment. Figure 1(c) shows the results of a dynamic difference-in-differences estimation, in which we find that the average estimated change in treated countries’ comments and signatures tends to be worse than the baseline week (Week -1) during the weeks post-AI access. For example, the top graph in Figure 1(c) shows that in the weeks after the full launch (first row, blue, right of Week 25), the share of petitions that received at least one comment within 30 days is consistently lower than the baseline week. During most weeks with AI access, the percent of petitions that reached one comment in 30 days declined by about 5 percentage points relative to the baseline week. Figure 1(d) shows that after AI tool release, the share of petitions that receive a comment in 30 days decreased by -1.11 percentage points relative to the share of petitions pre-AI tool release, although that difference was not statistically significant (95% CI = [-4.12, 1.90]). However, the share of petitions that reached 10 signatures when written with access to AI decreased by -4.65 percentage points, which was statistically significant (95% CI = [-7.74, -1.55]. In sum, we do not find evidence that petition outcomes are improved when users have access to AI.
In the SI Appendix we provide details of an exploratory predictive model showing that the lexical style of petitions written with access to AI (longer text, higher complexity) was associated that of petitions that had achieved positive outcomes prior to AI tool access. The predictive power of these features declined when users received access to the in-platform AI tool. We consider potential explanations and ramifications in the Discussion section.
Participation
(a) Word Count: Values and Means
(b) Percent of Petitions with 1+ Comment in 30 Days
Figure 3: Returning petition writers (N=4,553 users) from treated countries, on average, write longer petitions with worse outcomes when their second petition is written with access to AI. (a) Petition length is higher in petition writers’ second petitions, in all conditions. The greatest increase in word count occurs for writers in the Pre/Post split cohort. Average word count is shown in black with 95% CI’s. (b) Petition outcomes are worse in writers’ second petitions, in all conditions. The greatest drop in outcomes for users’ second petitions occurs when the second petition is written with access to AI. 95% CI’s (boostrapped, 1,000 iterations) are shown. Returning petition writers who return within 6 months are excluded, to enforce the same minimum time gap between petitions across cohorts.
Participation on Change.org did not change after users gained access to the in-platform AI tool. While the AI tool did not improve petition outcomes, one possible explanation could be that the tool’s launch attracted a different population. For example, the launch could have resulted in an influx of low-skill writers, thereby reducing the overall skill of writers on the platform. In this hypothetical case, platform-level trends might not improve, even if the tool improved the outcomes of individual petitions. However, we did not find evidence that the effect of AI access on platform trends was driven by changed participation patterns.
A time-series analysis of participation as measured by the volume of users writing petitions on the platform showed that participation did not increase with AI access. [^4] We were unable to conduct a differences-in-differences analysis of petition participation because pre-AI participation and volume trends varied between treated and control countries, violating the test requirements. Instead, the volume of user participation by week remained consistently in the same range as user participation prior to AI tool access.
To test the consistency of platform participation, we used pre-AI user participation data from treated and untreated countries to fit a Bayesian Structural Time Series (BSTS) model (see details in SI Appendix) [brodersen2015causalimpact]. This BSTS model uses pre-AI long-term trends and seasonal fluctuations to predict a synthetic control for counterfactual participation in treated countries had there been no AI tool launch. This forecast for platform participation closely reflected actual user participation, with actual participation falling within a 95% credible interval of forecast participation. Through comparing the synthetic counterfactual participation against actual platform participation, we do not find evidence of a significant effect of AI access on user participation, in either direction ($\Delta$=-0.45, 95% CI [-2.71, 1.75], p = 0.349).
Repeat Petition Writers
We further demonstrate that changes to platform composition were not responsible for changes to platform trends by showing that the platform-level trends—shifting petition features and unimproved outcomes—are consistent even when examining trends within repeat users. We focus our analysis here on one lexical feature (word count) and one outcome (comment engagement). The results for the other variables were similar, and are shown in the SI Appendix. We do not conduct a panel difference-in-differences analysis because of the rarity of repeat petition writing; only 7.6% of users in our dataset wrote multiple petitions, which amounted to too few untreated users for a reliable balanced panel.
We find that word count increases significantly between users’ first and second petitions when their first petition is written without AI tool access and their second petition is written with access to the AI tool (Pre/Post cohort); there is no significant change, however, when users have access to AI for neither of their petitions (Pre/Pre cohort) or both of their petitions (Post/Post cohort). Figure 3 shows word count (3(a)) and comment engagement (3(b)) in petitions written by three cohorts of returning writers: users who wrote both petitions without access to the AI tool (left facet of each figure), users who wrote their first petition without access and their second petition with access to the AI tool (middle), and users who wrote both petitions with access to the AI tool (right). Figure 3(a) shows that across cohorts, on average, returning writers’ second petitions were longer than their first petitions. For example, in the Pre/Post cohort (middle facet), writers’ first petitions were written without access to AI and had an average word count of 215 words (95% CI = [203.23, 227.40])(left column in orange). The same users’ second petitions were written with AI access, and had an average word count of 251 words (95% CI = [242.41, 259.34]) (right column in blue). We fit a linear mixed-effects regression model to predict log-transformed word count with fixed effects for cohort (Pre/Pre, Pre/Post, Post/Post) and petition order, and random effects for user ID. We then used pairwise contrasts with Bonferroni correction [scottpairwise1980] for multiple comparisons to test the contrasts between first and second petitions written within each cohort. Pairwise contrasts demonstrated that within the Pre/Post group, users’ second petitions were estimated to be 32.6% longer than their first petitions (ratio = 1.39, z = 20.183, p << .001). The estimated increase in first to second petition word count was smaller in the Pre/Pre group (ratio = 1.06, z = 2.459, p = 0.0418) and Post/Post group (ratio = 1.06, z = 2.108, p = 0.105). Overall, within repeat petition writers, the positive relationship between AI access and word count reaffirms the global trends we observed in the platform difference-in-differences analysis.
We also found that across cohorts, returning writers’ second petitions had worse outcomes than their first petitions, and that this decline was greater when users had AI access for their second petition. We use our preregistered primary outcome metric, the presence of at least one comment in 30 days threshold, as an indicator of petition outcomes. We saw similar results for the other outcomes, included along with model summaries in the SI Appendix. In the Pre/Post cohort (middle facet of 3(b)), 30% (95% CI = [28.35, 31.91]) of writers’ first petitions (first bar, orange) and 25% (95% CI = [23.77, 26.91]) of writers’ second petitions (second bar, blue) received at least one comment in 30 days. We estimated the predictive relationship of AI access on comment engagement by fitting a mixed-effects logistic regression model. Through Bonferroni-adjusted pairwise contrasts, we find that Pre/Post users’ second petitions, which were written with access to AI, have 34% lower odds of a positive outcome than their first petitions, with this difference being highly significant (OR=0.709, z=-4.61, $p<0.001$). On the other hand, the first and second petitions written by users in the Pre/Pre cohort, for whom both petitions were written without AI access, have no significant difference in outcomes (OR=0.871, z=-1.23, p=0.652). In the Post/Post cohort we observe a similar pattern to the Pre/Post cohort, with a significant difference in odds of first and second petition positive outcome (OR=0.595, z=-3.851, $p<0.001$). The decline in petition outcomes for petition writers’ second petitions across cohorts reflects a predictable regression to the mean. However, we also see that among repeat petition writers, AI access is associated with significantly worse petition outcomes. The negative relationship between AI access and petition outcomes among repeat-petition writers extends our findings from the platform-level difference-in-differences analysis, in which we see no evidence that AI access improved petition outcomes.
Finally, our repeat-writer data allowed us to see whether AI access was associated with improved outcomes for returning users whose first petitions had poor outcomes. Previous research found that AI has the greatest benefit for low-performing users [dellacquaNavigatingJaggedTechnological2023, Noy2023-sf]. We thus tested the relationship between cohort (Pre/Pre, Pre/Post, Post/Post) and second-petition outcomes among users who had unsuccessful first petitions. We again focus on petition outcomes through an indicator of whether petition writers receive at least one comment in 30 days. Of users who had unsuccessful first petitions (no comments within 30 days), we found no evidence that cohorts with AI access had better second-petition outcomes than expected compared to the other cohorts ($\chi^{2}=0.68,\text{df}=2,p=0.7115,N=3,147,w=0.015$). In summary, AI access was not associated with improved outcomes for low-performing users.
Discussion
Our work provides a causal analysis of the impact of AI integration into a real-world online platform. Our findings extend prior work on the impact of AI on online platforms at large: while online platforms might be influenced by both off-platform AI and in-platform AI tools, past studies have focused on general-use off-platform AI. As online platforms increasingly integrate AI, it is important to understand the potential impact on platform dynamics. In our case, the staggered roll-out of the Change.org in-platform AI tool allowed us to circumvent the challenges of noisy AI detection, and instead directly measure the impact of in-platform AI access through causal methods. We find that in-platform AI on Change.org shifts petition lexical features and increases site-wide homogeneity. At the same time, the in-platform AI tool did not improve platform outcomes, nor did it seem to affect platform participation. Over time, the in-platform AI integration may create undesirable new harms, such as increased suspicion by content consumers, or reduced effort by content producers.
Why doesn’t AI improve outcomes on Change.org? We have three hypotheses for the lack of change in outcomes: first, that AI might not substantively improve content quality; second, that AI might increase readers’ AI suspicion; and third, that AI-assisted writing might reduce petition-writer ownership and effort. Previous research suggests that when users adopt AI writing tools, their work scores higher in persuasion, creativity, and task performance [schoenegger2025large, Noy2023-sf, dellacquaNavigatingJaggedTechnological2023]. However, AI tools are not guaranteed to augment human performance. In a 2024 meta-review of human-AI collaboration, 15% of reviewed human-AI systems performed worse than humans alone. Dell’Acqua et al. explain the varied success of human-AI collaboration through the ever-expanding “jagged technological frontier” of AI capabilities [dellacquaNavigatingJaggedTechnological2023]. Within this frontier, AI augments human productivity and performance; outside of the frontier, AI performs worse than humans alone. As AI is capable of generating text that contains elements that have been shown to drive petition success, such as persuasion and emotional appeal [HAGEN2016783, chen2019multi, schoenegger2025large], we would expect AI to be capable of generating petitions that achieve strong outcomes. However, the technological capabilities of AI might be insufficient to explain the outcomes of content produced on social platforms, as evaluations are determined by readers’ dynamic, subjective, and social perceptions. Specifically, it is possible that petitions, with their real-world, context-specific missions, fall outside of AI’s capabilities for augmentation because the task requires personal and community-specific knowledge. AI on Change.org might contribute to what others have called “semantic garbage” or “AI slop” [feuerriegel2023research, floridi2020gpt, moller2025impact, tullis2025sifting]: although AI shapes petitions to resemble previous articulate, well-researched petitions, the changes may be superficial, and thus be perceived as less informative or authentic than human-written petitions.
Another way in which readers’ perceptions of petitions might change with in-platform AI access could be AI suspicion. Prior work shows that while people use biased heuristics to detect AI [jakesch2023human], when people suspect AI use, they may perceive the content as less trustworthy and assign lower evaluations to the writers [kadoma2025generative, reif2025socialpenalty, jakesch2019ai, hohenstein2023artificial]. If the AI-written petitions induce higher reader suspicion, outcomes are indeed likely to be affected. This hypothesis is difficult to test with our data since human perceptions of AI assistance may have shifted considerably in the year since our study’s petitions were posted [duan2024understanding, matatov2024examining]. On the other hand, there are reasons to discount the “AI suspicion” hypothesis in this case. Petition readers were not primed to suspect AI use on Change.org; during the ten-week period for which we estimate the causal impact of AI, Change.org provided no information to readers that petitions could have been written with AI. However, if AI suspicion is driving petition outcomes, this problem might worsen as Change.org increases their advertising, and consequently readers’ suspicion, of the AI tool.
Our third hypothesis for why petition outcomes were unchanged is not about how readers perceive others’ petitions, but how authors feel and act on their own petitions. Petition writers may feel diminished ownership over petitions they write with AI, and therefore be less likely to take actions that would attract signatures and comments, such as sharing the petition on social media. Previous literature on ownership in AI systems supports this explanation, showing that writers do not consider themselves to be the owners of text that is written with substantial AI support, especially when they use AI tools that require minimal user involvement [draxler2024ai, kadma2024inclusion, xu2024what, he2025which]. As perceptions of ownership contribute to how people construct their personal and collective identities [pierce2003state], and identity is central to sharing behaviors on social media [zhao2013faces], we would expect the diminished ownership to result in writers who are less likely to promote their petition in online communities. One way to address this concern could be through adding friction to the AI-assisted petition writing process, such as by requiring a lengthy user input or prompting the user to edit the draft produced by the AI tool. Research in platform design has demonstrated that friction can improve decision-making in collaborative human-AI systems by forcing users to have cognitive engagement [buccinca2021trust]. Such mechanisms could potentially improve authors’ perceptions of ownership and their likelihood of sharing petitions.
While petition outcomes remain the same, or potentially worsen, our data also indicates that Change.org participation is unchanged by AI access. Previous work has shown that general-purpose AI can cause participation to significantly grow on content sharing platforms [tullis2025sifting]. In particular, researchers, journalists, and community moderators have warned that the lowered cost of content production will enable the mass production of AI-generated misinformation and AI slop [feuerriegel2023research, floridi2020gpt, moller2025impact, tullis2025sifting, lloyd2023there]. However, if users only enter the petition-writing flow with strong intent to write a petition, the marginal increase in production could be limited, like we have seen in this case. Furthermore, since Change.org, unlike other platforms, does not offer financial or social incentives for mass production of content, users may have little incentive to create superfluous petitions. Overall, the participation trends offer a mixed bag for Change.org: on one hand, participation did not grow, despite prior research suggesting that AI assistance would broaden access to the platform. On the other hand, Change.org so far has likely avoided the fate of other platforms which are subjected to excessive AI spam and slop [tullis2025sifting].
It is especially important to consider the impact of AI integration on Change.org, where AI integration can influence which causes succeed and shift how users connect to their own communities. We found that AI has already caused a dramatic shift to petition lexical features on Change.org and driven platform-level homogenization. If AI continues to homogenize the language of protest, it has the capacity to undermine signals of petition authenticity and legitimacy [wojtowicz2025undermining]. Already, lexical features that were once highly predictive of petition outcomes have become weakened by platform-wide AI-driven shifts in lexical features. Regardless of lexical style, the introduction of in-platform AI can impact readers’ evaluations and writers’ engagement with their communities. Furthermore, with AI minimizing the cognitive load of petition-writing, petition writers might not only feel diminished ownership but also less engagement with the petition’s cause and their community at large. In sum, AI integration in petition-writing might significantly change the dynamics of civic engagement in ways that are difficult to predict and measure. Overall, our study serves as a cautionary tale for platforms that seek to integrate AI into their platforms: while user-generated content may change, AI tools may not yield dividends in participation or effectiveness, while potentially introducing additional harms.
Materials and Methods
Data
Our data consisted of 1,517,433 petitions posted to Change.org between January 1, 2022 and December 31, 2024. Change.org is the largest online petition service, providing a platform for individuals to create their own petitions or support others’ petitions. When users visit a Change.org petition, they are encouraged to sign the petition by sharing their name and zipcode. Their signature then counts toward the total count of signatures displayed on the petition’s page. Signees are then encouraged to support the petition by posting a comment. Comments are therefore only posted by petition supporters who already signed a petition, and can be considered an exceptionally strong vote of support for the petition. Petition writers can declare their petitions to be a victory at any point — however, victory status is a scarce and ambiguous signal. Some petitions are declared victories without ever accruing any signatures, and of the petitions we collected, only 1.38% were ever declared victories.
To collect the petitions posted to Change.org between January, 2022 and December, 2024 we used a web-scraping library [^5] https://docs.scrapy.org/en/latest/ to extract the text, comments, and metadata of all petitions posted during the time period. This dataset contains all petitions that had not been deleted or taken down as of the time of collection (January 9, 2025). Our data includes full petition text, title, and metadata about the author and the petition. Petition metadata includes user ID, engagement (signatures, comments), and labels such as language tag. Since each petition is connected to the author’s user ID, we could track all petitions written by the same user during the period for which we collected data. The main causal analysis of our study uses a subset (N = 277,496) of English language petitions that were posted by users in the United States, Canada, Great Britain, and Australia between January 1, 2022 and December 15, 2023666We subset to December 15 as Australia received in-platform AI access in late December. For robustness we repeat analysis with full month of December, in SI Appendix., and met minimum length requirements. Since we were interested in the lexical attributes of the text, this subset only included petitions with text that included at least two sentences and more than five words. We conducted a robustness check to estimate difference-in-differences results with the full text, including petitions with fewer than five words or two sentences. The SI Appendix contains details on summary statistics as well as additional robustness checks.
Petition Lexical Features
Our lexical features characterize petition text through length, lexical diversity, and reading grade level. We calculate these metrics on the combined text of the petition title and full text of the petition, with the title treated as a first sentence. Word count is measured as the number of tokens (separated by a space) in the title and full body of the petition.
Lexical diversity is calculated through the moving average type-token ratio (MATTR), which is resistant to varied text length [ZENKER2021100505]. The MATTR is calculated [^7] https://github.com/kristopherkyle/lexical_diversity by first taking the ratio of unique token types to total tokens for every window of a fixed size over the full text and title. In our analysis, each word (separated by a space) is a token, and unique token types are unique words. We use a window size of 50, as is typical in prior studies [ZENKER2021100505]. The MATTR of each petition is then the average of these ratios over all windows. Formally, the petition $p$ MATTR score is as follows, with $W_{p}$ as the set of the petition’s full text words, $N_{win}$ as the window size, and $U_{p,i}$ as the set of the petition’s unique words within each window of $N_{win}$ words starting with word $i$.
$\text{MATTR}(p)=\frac{1}{|W_{p}|-N_{win}+1}\sum_{i=0}^{|W_{p}|-N_{win}}\frac{|U_{p,i}|}{N_{win}}$ (1)
Readability is calculated through the Flesch-Kincaid grade level, for which higher grade levels correspond to more complicated text [kincaid1975derivation]. The Flesch-Kincaid grade level is a common metric to evaluate text complexity in a variety of high-impact contexts, such as education and public health [tanprasert2021flesch, jindal2017assessing]. The Flesch-Kincaid grade level metric seeks to capture complexity at both the sentence level, by considering average number of words per sentence, and word level, by considering average number of syllables per word. With the petition’s $p$ sets of words $W_{p}$, sentences $Se_{p}$, and syllables $Sy_{p}$, The Flesch-Kincaid grade level (FKGL) is calculated on the full text and title as follows [^8] https://github.com/cdimascio/py-readability-metrics:
$\text{FKGL}(p)=0.38\times\frac{|W_{p}|}{|Se_{p}|}+11.8\times\frac{|Sy_{p}|}{|W_{p}|}-15.59$ (2)
Petition Outcomes
On Change.org, signatures and comments are the core outcomes that drive a petition’s success. It is unlikely for a petition to have a positive outcome without broad support, as is demonstrated by signatures and comments. These indicators are the primary way by which petitions gain legitimacy, media attention, and ultimately convince decision-makers to enact change. We consider minimal thresholds of petition outcomes: that a petition must receive 10 signatures and have 1 comment in 30 days. It would be unlikely that any petition, regardless of scale or motivation, is able to reach any strong outcomes without achieving these minimum thresholds.
Comments are a signal of especially strong support from a signee, and demonstrate a higher level commitment for the petition. We report the number of comments written within the first 30 days after a petition is posted. We exclude comments written by the user who posted the petition. Even though most engagement occurs soon after a petition was posted, by limiting the time window to 30 days, we ensure that all petitions written during our period of analysis have the same opportunity for receiving a comment, and avoid biasing outcomes in favor of petitions that have been posted longer. While this was the primary outcome metric that we preregistered, our analysis shows that the time-boxing likely made no difference compared to using the full window. Of the 416,914 petitions that ever gain a comment and were posted at least 30 days before data collection, only 0.26% of petitions received their first comment outside of the 30-day window.
Signatures are the primary capital by which petitions gain legitimacy and are able to persuade decision-makers. The Change.org platform even states, “Petitions with 1,000+ supporters are 5x more likely to win!” We report petition signatures at the threshold of 10 total signatures. We do not have data on when petitions gain signatures. However, petitions included in our difference-in-differences analysis had between 1 and 3 years to gain signatures. Provided that less than 1% of petitions gained their first comment after the first 30 days, we expect that most petitions also reach 10 signatures within the first 30 days, if at all.
Difference-in-Differences Design
Change.org conducted a staggered roll-out of its AI writing tool by geographic region. The United States, Great Britain, and Canada were among the countries that participated in A/B testing during March-September 2023 and then received a full launch of the tool in October, 2023. However, there were several countries that had delayed access to the AI tool: Australia, New Zealand, and India. Other countries, like China, never received access. We used Australia as the control for our difference-in-differences analysis, as it showed adequately similar pre-AI trends in lexical features and outcomes in the United States, Great Britain, and Canada. While New Zealand, India, and China failed to demonstrate adequately similar pre-AI trends, we include petitions written in these countries in supplementary synthetic control analyses, included in the SI appendix.
We used static and dynamic difference-in-differences to measure the causal impact of access to the in-platform AI tool on petition lexical features and outcomes. For the static difference-in-differences approach, we use ordinary least squares (OLS) regression with country and week period fixed-effects, to control for differences in groups and over time. We measure time period fixed effects at the week level, following previous work, and use countries as the groups. We follow the difference-in-differences model specified by del Rio-Chanona et al. when studying the impact of AI on Stack Overflow participation [rio-chanonaAreLargeLanguage2024] where researchers measured production volume changes based on off-platform access to external LLMs such as ChatGPT.
$Y_{c,t}=\alpha_{c}+\lambda_{t}+\beta\cdot Treated_{c,t}+\boldsymbol{\eta}\cdot\mathbf{X_{c,t}}+\epsilon_{c,t}$ (3)
Here, the dependent variable $Y_{c,t}$ reflects the metric for which we are estimating change. When studying lexical features we use the median lexical feature (e.g. median word count) as the dependent variable. When studying outcomes, we use the share of petitions that achieved the outcome threshold (e.g. share of petitions reaching 10 signatures). $\alpha_{c}$ are country fixed effects, $\lambda_{t}$ are time-period (i.e. week) fixed effects, and $\epsilon_{c,t}$ is the error term. Pre-treatment covariates can be included through the model matrix $\mathbf{X_{c,t}}$ and coefficients of these covariates through $\boldsymbol{\eta}$. We do not include any covariates in our main analysis; the SI Appendix provides robustness checks with additional covariates, such as day of week and petition topic. The $Treated_{c,t}$ term reflects the interaction between time period and country, thereby taking the value of 1 if a country had AI access during a given week, and 0 otherwise. The value $\beta$ thus captures the average estimated treatment effect (ATT) of access to the in-platform AI tool in countries that received access. We cluster standard errors at the month level, again following previous difference-in-differences studies of online platform activity [rio-chanonaAreLargeLanguage2024].
In total, 277,496 petitions are considered for the difference-in-differences analysis, which includes 65 weeks pre-AI and 10 weeks post-AI, and excludes the AI tool’s experimental A/B test period as noted above. We create our static difference-in-difference estimation using the doubly robust method proposed by Sant’Anna and Zhao [SANTANNA2020101]. To show that our analysis is not specific to the time period, we repeated our analysis on a narrower pre-AI time frame and show that results hold (SI Appendix).
In addition to calculating aggregate estimates for the effect of AI in treated countries, we conducted a dynamic difference-in-differences analysis to measure how lexical features and outcomes changed dynamically over time periods. In this setting we create difference-in-difference estimates that are grouped by time period (week) relative to a universal baseline (the week prior to A/B testing). We used the week prior to A/B testing (Week -1) as a reference point to estimate the effect of varying lengths of exposure to the treatment (access to in-platform AI). To calculate the dynamic difference-in-differences estimates shown in Figures 1(a) and 1(c), we used the method and software developed by Sant’Anna and Callaway [callaway_santanna_2025_did] to measure group-time average treatment effects that account for heterogeneity [callaway2021difference]. In particular, instead of calculating the overall average treatment effect on treated users (ATT) that is captured in equation 3, Callaway and Sant’Anna calculate ATT(g,t) values, which are specific to units in group $g$ that are treated at time $t$, as shown in equation 4. Here, $Y_{t}(g)$ reflects the outcome of group $g$ at time t, and $Y_{t}(0)$ reflects the potential outcome of group $g$ at time $t$ had it never been treated. Since our causal analysis terminates before any of the delayed-release countries receive AI access, all countries are all either untreated or treated across the full time period, and as such our ATT(g,t) estimates simplify to ATT(t) estimates for each time period.
$ATT(g,t)\;=\;\mathbb{E}\!\left[\,Y_{t}(g)\;-\;Y_{t}(0)\;\middle|\;G=g\,\right]$ (4)
Repeat Petition Writers
For our study of repeat petition writers, we sought to understand how petition lexical features and outcomes differed between users’ first and second petitions, and how this difference varied by AI access.
We first fit a mixed-effects linear regression model to predict word count (log transformed due to the right skew) among repeat petition users. In this model we have fixed effects for user cohort (Pre/Pre, Pre/Post, Post/Post), petition order (First, Second), and the interaction between cohort and petition order. We also include random effects for user ID, as we have repeated measures for each user. The word count model is captured in Eq. 5.
$\log(\text{Word Count}_{ij})=\beta_{c}\cdot\mathbf{1}\{Cohort_{ij}=c\}\beta_{p}\cdot\mathbf{1}\{PetitionOrder_{ij}=p\}+$
$\beta_{c,p}\cdot(\mathbf{1}\{Cohort_{ij}=c\}\times\mathbf{1}\{PetitionOrder_{ij}=p\})\gamma_{i}+\epsilon_{ij}$ (5)
In this model, $j$ reflects the observations for each user ID, $i$. $\gamma_{i}$ reflects the random effects attributed to each user ID. We are primarily focused on the coefficient of the interaction between cohort and petition order, which is captured by $\beta_{c,p}$. We compare the levels of the interaction term $\beta_{c,p}$ through pairwise contrasts, in which each contrast corresponds to a difference in expected log-word count between the first and second petitions for a given cohort. We then adjust the pairwise contrast estimates with a Bonferroni correction, to account for multiple hypothesis testing [scottpairwise1980]. In the SI appendix we use the same model specification to understand the predictive relationship between repeat-users cohort and our other lexical measures, MATTR and Flesch-Kincaid grade level.
We fit a similar logistic mixed-effects model to predict differences in petition outcomes among repeat petition users. We use the same specification of random and fixed effects as in Eq. 5, but fit a logistic rather than linear regression model as our dependent variable, whether a petition received a comment in 30 days, is a binary value. The model is shown in Eq 6.
$\log\left(\frac{P(Outcome_{ij}=1)}{1-P(Outcome_{ij}=1)}\right)=\beta_{c}\cdot\mathbf{1}\{Cohort_{ij}=c\}+\beta_{p}\cdot\mathbf{1}\{PetitionOrder_{ij}=p\}+$
$\beta_{c,p}\cdot(\mathbf{1}\{Cohort_{ij}=c\}\times\mathbf{1}\{PetitionOrder_{ij}=p\})+\gamma_{i}$ (6)
We again compare pairwise contrasts, adjusted for multiple hypothesis testing, between levels of the interaction term. In this model, we interpret contrasts as explaining the difference in log-odds of a positive outcome between first and second petition order, depending on cohort. Through these pairwise contrasts, we can compare the association between petition order and outcome variable across user cohorts.
Data Availability
We make petition metadata available at the following OSF project link, included in footnote here. [^9] https://osf.io/q3hj5/?view_only=3fc5b89d05eb4741957af431b107cf2a All analyses included in the main text and figures, as well as in the supplementary material analysis and figures, can be reproduced with the data and code shared here. [^10] We do not make full petition text data or user names available, even though they are publicly available, for user agency, privacy, and copyright concerns. The full petition text of all non-deleted petitions can be retrieved from the Web using the petition IDs provided in the data. Published academic researchers will be able to request the full text dataset from the authors.
Supporting Information
si_main